The angle of a sector is 30deg and the radius is 15cm.




Find the area of the sector.




10. A circle having an area of 224sq.m is inscribed in an




octagon. Find the area of the octagon.




11. A circle is circumscribed about a hexagon. Determine




the area of the hexagon if the area outside the hexagon




but inside the circle is 15sq.cm.





1
Expert's answer
2021-11-16T19:15:32-0500

9) Area of a sector ="\frac{\theta }{360}\pi r^2=\frac{30}{360}\times \frac{22}{7}\times 15^2=58.93cm^2"

10) Area of the circle "=\pi r^2=224"

"r=\sqrt{224/\pi}=8.4444m"

From the centre of the octagon, it can be divided into 8 equal triangles.

Angle at the centre for one triangle "=360\/8=45^0"

Height of one triangle =radius of the circle =8.4444m

Base angles of the triangle "=\frac{180^0-45^0}{2}=67.5^0"

Hypotenuse of the triangle "=\frac{8.444}{sin67.5^0}=9.14m"

Area of the octagon =area of one triangle "\times 8=1/2\times 9.14 \times sin 45^0\times 8=236.29m^2"


11) A hexagon can be divided into 6 equal triangles

Angle at the centre "=360^0\/6=60^0"

The triangles are equilateral, meaning all sides are equal = x cm

The area of the hexagon if the area outside the hexagon but inside the circle"= (60^0/360^0\times \pi \times x^2-1/2 \times x^2 \times sin60^0)6=15"

"0.5236x^2-0.433x^2=15\/6=2.5"

"0.0906x^2=2.5"

"x^2=27.5938"

"x=5.253cm"

Area of the hexagon = area of one triangle "\times 6=1/2\times 5.253\times 5.253\times sin60^0\times 6=71.69cm^2"